18.-/4 points SerPSE9 5.P.028.WI 0 My Notes Ask Your Teacher The systems shown below are in equilibrium with m 6.30 kg and 32.0°. If the spring scales are calibrated in newtons, what do they read? Ignore the masses of the pulleys and strings and assume the pulleys and the incline are frictionless. Im In scale in (a):N scale in (b):- N scale in (c): N scale in (d):N Need Help? Watch N
Answer #1

in a)

reading = m * g =6.3 * 10

reading = 63 N

in b)

reading = tension = m * g =6.3 * 10

reading = 63 N

in c)

reading = 2m * g = 2 * 6.3 * 10

reading = 126 N

in d)

reading = m * g * sin(31 degree) = 6.6 * 10 * sin(32)

reading = 34 N